OpenStudy (anonymous):

-2/5x -9 < 9/10 Solve and show steps.

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

Help meee lol

OpenStudy (anonymous):

what do I do first

OpenStudy (anonymous):

This? \[\large -\frac{2}{5}x-9<\frac{9}{10}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, it might help if you multiply everything by 10, first.

OpenStudy (anonymous):

Get rid of those pesky denominators.

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

Well, you could multiply both sides by 10, as permitted by the Multiplication property of Inequality... \[\large 10\left(-\frac{2}{5}x-9\right)<10\left(\frac{9}{10}\right)\]

OpenStudy (anonymous):

You'll get \[\large -4x -90<9\] And just solve :)